OpenStudy (anonymous):

Convert 149° in Fahrenheit to Celsius temperature using the formula . 9/5c+32. PLEASE HELP

OpenStudy (anonymous):

149 = 9/5 C + 32 149 - 32 = 9/5 C + 32 - 32 117 = 9/5 C solve for C

OpenStudy (anonymous):

my answer choices are 70c 75c 60c 65c